OpenStudy (mirandaregina):

I need help using the substitution method to solve the system 2x + 5y = 10 -3x + y = 36

OpenStudy (anonymous):

Ok in the second equation how would you rewrite it to show as 5y =????

OpenStudy (mirandaregina):

@nelsonjedi would equation 2 be revised as y = 3x + 36 ?

OpenStudy (anonymous):

Yes. Now we need to reflect the y values as 5y so we can Substitute that value in the first equation. So take y=3x+36 would equate to 5y=???????

OpenStudy (mirandaregina):

I have: y = 3x + 36 2x + 5y = 10 sub revised eq. 2 into eq. 1 2(3x + 36) + 5y = 10 6x + 72 + 5y = 10 I'm not sure what comes next; this is where I need help

OpenStudy (anonymous):

Hi, here's a link I think will help a bit: http://www.webmath.com/solver2.html

OpenStudy (anonymous):

In substitution you are trying to eliminate one of the unknowns (x or y). So if we take y=3x+36 and equate it to 5y...we would multiply each component by 5 and arrive at 5y=15x+180...Do you see this?

OpenStudy (mirandaregina):

So, y = 3x + 36 5y = 15x + 180 ?

OpenStudy (mirandaregina):

@Matty13579 thank you :)

OpenStudy (anonymous):

Yes..now you have a value for 5y from the 2nd equation plug that value into the first equation (substitute) and you can solve for x.

OpenStudy (mirandaregina):

2x + 5y = 10 2x + 5(15x + 180) = 10 2x + 72 + 900 = 10 74x = 10 + 900 (I'm not sure if this is what I'm supposed to do here)

OpenStudy (anonymous):

5y =15x+180 so you do not have to multiply it by 5. When you substitute you will have 2x+15x+180=10 Then combine terms to get 17x+180=10. Then subtract 180 from both sides and you get 17x= -170...so what does x equal?

OpenStudy (mirandaregina):

17x = -170 x = -10

OpenStudy (anonymous):

Now that you know x you can solve for the value of y by plugging it into either of the equations

OpenStudy (mirandaregina):

y = 3(-10) + 36 y = -30 + 36 y = 6

OpenStudy (anonymous):

@Mirandaregina You're welcome!

OpenStudy (mirandaregina):

@nelsonjedi thank you very much for helping me out with this :)

OpenStudy (texaschic101):

Miranda...the only thing you did wrong with your problem is that you subbed in 3x + 36 in for the x value and not the y value. Your problem should have been set up like this : y = 3x + 36 2x + 5y = 10 2x + 5(3x + 36) = 10 and just solve for x from here

OpenStudy (mirandaregina):

@texaschic101 ah ok, that's very helpful to know. thank you!